### v1

@@ Abstract
Orthogonal matching pursuit (OMP) is a commonly used greedy algorithm for recovering sparse signals from compressed measurements. In this paper, we introduce a variant of the OMP algorithm to reduce the complexity of reconstructing a class of $K$-sparse signals $\boldsymbol{x}\in\mathbb{R}^{n}$ from measurements $\boldsymbol{y}=\boldsymbol{A}\boldsymbol{x}$, where $\boldsymbol{A}\in\{0,1\}^{m\times n}$ is a sparse random combinatorial matrix with $d~{}(d\leq m/2)$ ones per column. The proposed algorithm, referred to as the confined OMP algorithm, utilizes the properties of $\boldsymbol{x}$ and $\boldsymbol{A}$ to remove much of the redundancy in the dictionary (also referred to as $\boldsymbol{A}$) and thus fewer column indices of $\boldsymbol{A}$ need to be identified. To this end, we first define a confined set $\Gamma$ with $|\Gamma|\leq n$ and then prove that the support of $\boldsymbol{x}$ is a subset of $\Gamma$ with probability 1 if the distributions of non-zero components of $\boldsymbol{x}$ satisfy a certain condition. During the process of the confined OMP algorithm, the possibly chosen column indices are strictly confined into the confined set $\Gamma$. We further develop lower bounds on the probability of exact recovery of $\boldsymbol{x}$ using OMP algorithm and confined OMP algorithm with $K$ iterations, respectively. The obtained theoretical results of confined OMP algorithm can be used to optimize the column degree $d$ of $\boldsymbol{A}$. Finally, experimental results show that the confined OMP algorithm is more efficient in reconstructing a class of sparse signals compared to the OMP algorithm.
keywords:
Compressed sensing, exact recovery probability, orthogonal matching pursuit, sparse signal recovery, sparse random combinatorial matrices
1 Introduction
@@
Compressed sensing (CS) as a novel sampling theory [6, 12] has attracted much attention in recent twenty years. In CS, it is common to encounter the following linear model
(1) $\boldsymbol{y}=\boldsymbol{A}\boldsymbol{x},$
@@
where $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ is a measurement matrix with $m<n$, $\boldsymbol{x}\in\mathbb{R}^{n}$ is an unknown $K$-sparse signal (i.e., $\boldsymbol{x}$ has at most $K$ non-zero elements) and $\boldsymbol{y}\in\mathbb{R}^{m}$ is a known measurement vector. The CS recovery algorithm aims to recover $\boldsymbol{x}$ from Eq. 1 by solving the following minimization problem
(2) $\min\lVert\boldsymbol{x}\rVert_{0}~{}~{}\mathrm{s.t.}~{}\boldsymbol{y}=\boldsymbol{A}\boldsymbol{x},$
@@
where $\lVert\boldsymbol{x}\rVert_{0}\overset{\triangle}{=}|\{i:x_{i}\neq 0\}|$ and $x_{i}$ is the $i$-th element of $\boldsymbol{x}$. Unfortunately, as shown in [18], the problem Eq. 2 is NP-hard in general. Two methods are commonly used for solving this problem. One focuses on the reconstruction of sparse signals by considering a convex relaxation of Eq. 2, such as solving the $l_{1}$-minimization problem [5, 36, 7, 15]. There are also algorithms solving Eq. 2 directly, such as greedy algorithms [35, 30, 8, 29, 32, 25, 37, 24, 39] and thresholding algorithms [4, 17, 33, 11, 2].
Among greedy algorithms, the OMP algorithm [35] is one of the most commonly used algorithms. As a greedy algorithm, the OMP algorithm identifies the support (index set of non-zero elements) of the sparse signal $\boldsymbol{x}$ in an iterative manner and thus iteratively performs local optimal updates. Specifically, the process of the OMP algorithm at each iteration can be summarized into four steps [37] (see Algorithm 1 for details):
Among these, the computational complexity of the OMP algorithm is mainly dominated by the identification step and the estimation step. In order to enhance the computational efficiency and recovery performance of the OMP algorithm, there have been some studies on the modified OMP algorithm, mainly focusing on the identification step. For example, the generalized OMP algorithm [37] (a.k.a. orthogonal multi-matching pursuit algorithm [25]) allows multiple indices maximally correlated with the residual are chosen at each iteration so that fewer number of iterations are required. The existing methods seek to efficiently select the “true” column indices from the redundant dictionary (measurement matrix $\boldsymbol{A}$). This raises a question: Is it possible to reduce the redundancy of the dictionary $\boldsymbol{A}$?
@@
  Input: $\boldsymbol{y}\in\mathbb{R}^{m}$, $\boldsymbol{A}\in\mathbb{R}^{m\times n}$, and $K$.
@@
     $t^{(k)}=\underset{1\leq i\leq n}{\operatorname*{arg\,max}}|\boldsymbol{A}_{i}^{T}\boldsymbol{r}^{(k-1)}|$(Identification)
@@
Before answering this question, we introduce some useful tools to characterize the performance of the recovery algorithm. In [6], Candès and Tao introduced the concept of restricted isometry property (RIP) and showed that if $\boldsymbol{A}$ satisfies RIP with relatively small restricted isometry constant (RIC) $\delta_{2K}$, any $K$-sparse signal can be exactly recovered by solving a $l_{1}$-minimization problem. In particular, it has been proved in [41] that $\delta_{K+1}<1/\sqrt{K+1}$ is sufficient for the OMP algorithm to recover any $K$-sparse signal from Eq. 1 in $K$ iterations. The mutual coherence, denoted as $\mu_{m}$, is also an important parameter for $\boldsymbol{A}$, which indicates the maximum absolute correlation between normalized columns of $\boldsymbol{A}$. It has been showed in [34] that any $K$-sparse signal can be exactly recovered by the OMP algorithm if $K<(\mu_{m}^{-1}+1)/2$. In [35], the authors developed a lower bound on the probability that any $K$-sparse signal can be exactly recovered from Eq. 1 by using the OMP algorithm in $K$ iterations, where the matrix $\boldsymbol{A}$ in Eq. 1 is a random Gaussian matrix. In [40], this lower bound is further tightened with the aid of prior information of $\boldsymbol{x}$. Unfortunately, the lower bound techniques proposed in [35] and [40] are only suitable for the case of Gaussian matrices.
@@
In CS, the construction of measurement matrices is also one of the main concerns. In general, the random measurement matrices can be classified into dense and sparse matrices. It has been verified that many dense matrices, such as Gaussian matrices and Fourier matrices, satisfy the RIP with overwhelming probability [1] and have provably good recovery performance. On the other hand, sparse random matrices also attract much attention [3, 22, 23, 10, 26, 28] since the sparsity can enable the computation of the matrix-product to be remarkably efficient and save the storage space in practice [19]. In particular, sparse binary-valued measurement matrices are commonly used in some applications, including group testing [20], DNA Microarrays [31], and single-pixel imaging [14]. Furthermore, many studies [10, 26, 27] showed that the sparse binary-valued measurement matrices are as “good” as the dense ones both in theory and in practice.
@@
Assume that the measurement matrix $\boldsymbol{A}\in\{0,1\}^{m\times n}$ is a sparse combinatorial matrix with independent columns, where each column is chosen uniformly among the vectors with $d~{}(d\leq m/2)$ ones. Now we answer the previous question. Yes, much redundancy of $\boldsymbol{A}$ can be removed if $\boldsymbol{x}$ is a signal defined in Definition 2.7. Actually, this observation can be traced back to the work of Khajehnejad et al. in [23]. They showed that the redundancy of a sparse matrix constructed by the expander theory can be eliminated if non-zero elements of $\boldsymbol{x}$ are non-negative. Our work is more general and takes the observation in [23] as a special case. Specifically, the sparse signal we consider is not limited to be non-negative, but a class of signals defined in Definition 2.7, including the Gaussian sparse signal. Furthermore, the considered measurement matrix $\boldsymbol{A}$ in this paper is more general than that constructed by the expander theory in [23]. The contributions of this paper are summarized as follows.
@@
  • We first define the confined set $\Gamma$ with $|\Gamma|\leq n$ and prove that the support of $\boldsymbol{x}$ defined in Definition 2.7 is a subset of $\Gamma$ with probability 1. To theoretically clarify the effectiveness of removing the redundancy of $\boldsymbol{A}$, we present the expectations of the sparsity of $\boldsymbol{y}$ and the size of $\Gamma$.
@@
  • We propose a variant of OMP algorithm, referred to as the confined OMP algorithm, by introducing the confined set $\Gamma$ into the identification step. The possibly chosen column indices are strictly confined into the confined set $\Gamma$. We further analyze the complexities of OMP algorithm and confined OMP algorithm. The analysis results show that the identification efficiency of confined OMP algorithm is at least $\frac{nKd-K}{|\Gamma|Kd-K+n}$ times that of OMP algorithm. Furthermore, the experimental results show that the confined OMP algorithm achieves a large reduction in complexity if $K\ll m$.
  • We develop a lower bound on the probability of exact recovery of any $K$-sparse signal $\boldsymbol{x}$ using OMP algorithm over a sparse random combinatorial matrix. As far as we know, our work is the first to develop the recovery lower bound of OMP algorithm over a sparse random combinatorial matrix. We further develop a lower bound on the probability of exact recovery of $\boldsymbol{x}$ defined in Definition 2.7 using confined OMP algorithm over a sparse random combinatorial matrix.
@@
The paper is organized as follows. We define the confined set $\Gamma$ and present the proposed algorithm in Section 2. In Section 3, the expectations of the sparsity of $\boldsymbol{y}$ and the size of $\Gamma$ are investigated, and then the recovery performance bounds of OMP algorithm and confined OMP algorithm are provided. Experimental results are presented in Section 4 and Section 5 concludes the paper.
@@
Notation: we use boldface lowercase letters to denote column vectors and boldface uppercase letters to denote matrices. The support of $\boldsymbol{x}$ is denoted by $\Omega$ and the complement of $\Omega$ is $\Omega^{c}=[n]\backslash\Omega=\{i:i\in[n],i\notin\Omega\}$, where $[n]$ represents the set $\{1,2,\cdots,n\}$. Let $x_{i}$ and $\boldsymbol{A}_{j}$ be the $i$-th element of $\boldsymbol{x}$ and the $j$-th column of $\boldsymbol{A}$, respectively. We use $A_{i,j}$ to denote the element of $\boldsymbol{A}$ located at the $i$-th row and the $j$-th column. We denote by $\boldsymbol{x}_{\Lambda}$ the sub-vector of $\boldsymbol{x}$ that contains the entries of $\boldsymbol{x}$ indexed by the set $\Lambda$, and $\boldsymbol{A}_{\Lambda}$ the sub-matrix of $\boldsymbol{A}$ that contains the columns of $\boldsymbol{A}$ indexed by the set $\Lambda$. We use $\boldsymbol{A}^{T}$ to denote the transport of $\boldsymbol{A}$ and $\mathbf{E}[X]$ to represent the expectation of $X$.
2 Proposed algorithm
@@
This section will first introduce the inherent properties between random combinatorial matrices and sparse signals, and then present the details of the proposed algorithm.
2.1 The confined set
@@
For a sufficiently small number $\epsilon>0$, let =$\mathcal{E}=\{i:|y_{i}|\leq\epsilon\}$ where $y_{i}$ is the $i$-th element of $\boldsymbol{y}$. The definition of the confined set is as follows.
Definition 2.1 (The confined set).
@@
Theorem 2.2 (The probability of $\{\Omega\subseteq\Gamma\}$).
Suppose that in Eq. 1, $\boldsymbol{A}\in\{0,1\}^{m\times n}$ is a random combinatorial matrix with independent columns, where each column is chosen uniformly among the vectors with $d$ ones. Furthermore, the $K$ non-zero components of $\boldsymbol{x}$ are independent and identically distributed (i.i.d.), with the same cumulative distribution function (CDF) $F_{X}(x)=\mathbb{P}\{X\leq x\}$. Then, the probability of $\{\Omega\subseteq\Gamma\}$ is given by
(3) $\mathbb{P}\{\Omega\subseteq\Gamma\}=\left(1-\sum_{\ell=1}^{K}\binom{K}{\ell}\Big{(}F_{X}^{*\ell}(\epsilon)-F_{X}^{*\ell}(-\epsilon)\Big{)}\right)^{|\mathcal{E}|}.$
@@
Proof 2.3.
See Appendix A.
With Theorem 2.2, we have the following two Corollaries.
Corollary 2.4.
If the probability density function (PDF) $f_{X}(x)$ of non-zero components of $\boldsymbol{x}$ is a continuous function, then the probability $\mathbb{P}\{\Omega\subseteq\Gamma\}\to 1$ as $\epsilon\to 0$.
Proof 2.5.
Obviously, $F_{X}^{*\ell}(x)$ is a continuous function provided that $f_{X}(x)$ is a continuous function. According to Theorem 2.2, we have $\underset{\epsilon\to 0}{\lim}\mathbb{P}\{\Omega\subseteq\Gamma\}=1$.
Corollary 2.6.
If the values of non-zero components of $\boldsymbol{x}$ share the same polarity, then the probability $\mathbb{P}\{\Omega\subseteq\Gamma\}\to 1$ as $\epsilon\to 0$.
The Corollaries indicate that by setting $\epsilon\to 0$, both Gaussian sparse signals and non-negative sparse signals can ensure $\mathbb{P}\{\Omega\subseteq\Gamma\}=1$. We call these signals as confined signals in this paper. The definition of the confined signal is given as follows.
Definition 2.7 (The confined signal).
The confined signal is a $K$-sparse signal whose non-zero components are i.i.d. and, for $\ell=1,2,\cdots,K$, their CDFs satisfy $F_{X}^{*\ell}(\epsilon)-F_{X}^{*\ell}(-\epsilon)\to 0$ as $\epsilon\to 0$.
2.2 Confined OMP algorithm
The confined OMP is a modified OMP algorithm by introducing a confined set $\Gamma$. The key feature of the confined OMP algorithm is to introduce a confined set $\Gamma$ into the identification step such that the possibly chosen column indices are strictly confined into the confined set $\Gamma$. In the following, we assume that the signal $\boldsymbol{x}$ is defined in Definition 2.7 and has exactly $K$ non-zero elements. The details of the confined OMP algorithm are summarized in Algorithm 2. The necessary explanations of the confined OMP algorithm are as follows.
  • (Preprocessing) For a sufficiently small number $\epsilon$, obtain the set =$\mathcal{E}=\{i:|y_{i}|\leq\epsilon\}$ first. Then, obtain the confined set $\Gamma=[n]\backslash\bigcup_{i\in\mathcal{E}}\Gamma_{i}^{c}$, where $\Gamma_{i}^{c}=\{j:A_{i,j}=1,~{}|y_{i}|\leq\epsilon\}$. With $\Omega\subseteq\Gamma$, we have $\Omega=\Gamma$ if they have the same size. Thus, the identification is already done without the subsequent iterations.

### v2

@@ Abstract
Orthogonal matching pursuit (OMP) is a commonly used greedy algorithm for recovering sparse signals from compressed measurements. In this paper, we introduce a variant of the OMP algorithm to reduce the complexity of reconstructing a class of $K$-sparse signals $\boldsymbol{x}\in\mathbb{R}^{n}$ from measurements $\boldsymbol{y}=\boldsymbol{A}\boldsymbol{x}$. In particular, $\boldsymbol{A}\in\{0,1\}^{m\times n}$ is a sparse random combinatorial matrix with independent columns, where each column is chosen uniformly among the vectors with exactly $d~(d\leq m/2)$ ones. The proposed algorithm, referred to as the confined OMP algorithm, leverages the properties of the sparse signal $\boldsymbol{x}$ and the measurement matrix $\boldsymbol{A}$ to reduce redundancy in $\boldsymbol{A}$, thereby requiring fewer column indices to be identified. To this end, we first define a confined set $\Gamma$ (see Definition 1 in Sec. II-A) with $|\Gamma|\leq n$ and then prove that the support of $\boldsymbol{x}$ is a subset of $\Gamma$ with probability 1 if the distributions of nonzero components of $\boldsymbol{x}$ satisfy a certain condition. During the process of the confined OMP algorithm, the possibly chosen column indices are strictly confined to the confined set $\Gamma$. We further develop the lower bound on the probability of exact recovery of $\boldsymbol{x}$ using the confined OMP algorithm. The derived bound shows that, if the spark of $\boldsymbol{A}$ is greater than $K$, then $m=2\mathrm{e}K\ln(n-K)$ measurements are sufficient to ensure the probability of recovering a $K$-sparse signal with the proposed algorithm is at least $1-\frac{1}{n-K}$, where $\mathrm{e}$ represents Euler’s number. Furthermore, the obtained theoretical results can be used to optimize the column degree $d$ of $\boldsymbol{A}$. Finally, experimental results show that the confined OMP algorithm is more efficient in reconstructing a class of sparse signals compared to the OMP algorithm.
I Introduction
@@
Compressed sensing (CS) as a novel sampling theory [1, 2] has attracted much attention over the past twenty years. In CS, it is common to encounter the following linear model
$\boldsymbol{y}=\boldsymbol{A}\boldsymbol{x},$ (1)
@@
where $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ is a measurement matrix with $m<n$, $\boldsymbol{x}\in\mathbb{R}^{n}$ is an unknown $K$-sparse signal (i.e., $\boldsymbol{x}$ has at most $K$ nonzero elements) and $\boldsymbol{y}\in\mathbb{R}^{m}$ is a known measurement vector. The CS recovery algorithm aims to recover $\boldsymbol{x}$ from (1) by solving the following $l_{0}$-minimization problem
$\min\lVert\boldsymbol{x}\rVert_{0}~~\mathrm{s.t.}~\boldsymbol{y}=\boldsymbol{A}\boldsymbol{x},$ (2)
@@
where the $l_{0}$ norm $\lVert\boldsymbol{x}\rVert_{0}$ is the number of nonzero elements in $\boldsymbol{x}$, i.e., $\lVert\boldsymbol{x}\rVert_{0}\overset{\triangle}{=}|\{i:x_{i}\neq 0\}|$. Here, $x_{i}$ is the $i$-th element of $\boldsymbol{x}$ and $|\mathcal{S}|$ is the cardinality of a set $\mathcal{S}$. Unfortunately, as shown in [3], the problem (2) is NP-hard in general. Two methods are commonly used for solving this problem. One focuses on the reconstruction of sparse signals by considering a convex relaxation of (2), such as solving the $l_{1}$-minimization problem [4, 5, 6, 7]. There are also algorithms solving (2) directly, such as greedy algorithms [8, 9, 10, 11, 12, 13, 14, 15, 16] and thresholding algorithms [17, 18, 19, 20, 21].
@@
 Input: $\boldsymbol{y}\in\mathbb{R}^{m}$, $\boldsymbol{A}\in\mathbb{R}^{m\times n}$, $\varepsilon$, and $K$.
@@
  $t^{(k)}=\underset{i\in[n]\backslash\Lambda^{(k-1)}}{\operatorname*{arg\,max}}|\boldsymbol{A}_{i}^{T}\boldsymbol{r}^{(k-1)}|$(Identification)
@@
Among greedy algorithms, the OMP algorithm [8] is one of the most commonly used algorithms. As a greedy algorithm, the OMP algorithm identifies the support (index set of nonzero elements) of the sparse signal $\boldsymbol{x}$ in an iterative manner and thus iteratively performs local optimal updates. Specifically, the process of the OMP algorithm at each iteration can be summarized into four steps [14] (see Algorithm 1 for details):
  • Stopping criteria: the iteration process is terminated when the number of iterations reaches the maximum value $K$ or the $l_{2}$ norm of the residual $\boldsymbol{r}^{(k)}$ falls below a preset threshold $\varepsilon$.
Among these, the computational complexity of the OMP algorithm is mainly dominated by the identification step and the estimation step. In order to enhance the computational efficiency and recovery performance of the OMP algorithm, there have been some studies on the modified OMP algorithm, mainly focusing on the identification step. For example, the generalized OMP algorithm [14] (a.k.a. orthogonal multi-matching pursuit algorithm [13]) allows multiple indices maximally correlated with the residual to be chosen at each iteration so that fewer number of iterations are required. The existing methods seek to efficiently select the “true” column indices from the redundant dictionary (measurement matrix $\boldsymbol{A}$). This raises a question: Is it possible to reduce the redundancy of the dictionary $\boldsymbol{A}$?
@@
Before answering this question, we introduce some useful tools to characterize the performance of the recovery algorithm. In [1], Candès and Tao introduced the concept of restricted isometry property (RIP) and showed that if $\boldsymbol{A}$ satisfies RIP with relatively small restricted isometry constant (RIC) $\delta_{2K}$, any $K$-sparse signal can be exactly recovered by solving a $l_{1}$-minimization problem. In particular, it has been proved in [22] that $\delta_{K+1}<1/\sqrt{K+1}$ is sufficient for the OMP algorithm to recover any $K$-sparse signal from (1) in $K$ iterations. The mutual coherence, denoted as $\mu_{m}$, is also an important parameter for $\boldsymbol{A}$, which indicates the maximum absolute correlation between normalized columns of $\boldsymbol{A}$. It has been shown in [23] that any $K$-sparse signal can be exactly recovered by the OMP algorithm if $K<(\mu_{m}^{-1}+1)/2$. In [8], the authors developed a lower bound on the probability that any $K$-sparse signal can be exactly recovered from (1) by using the OMP algorithm in $K$ iterations, where the matrix $\boldsymbol{A}$ in (1) is a random Gaussian matrix. In [24], this lower bound is further tightened with the aid of prior information of $\boldsymbol{x}$. Unfortunately, the lower bound techniques proposed in [8] and [24] are only suitable for the case of Gaussian matrices.
@@
In CS, the construction of measurement matrices is also one of the main concerns. In general, the random measurement matrices can be classified into dense and sparse matrices. It has been verified that many dense matrices, such as Gaussian matrices and Fourier matrices, satisfy the RIP with overwhelming probability [25] and have provably good recovery performance. On the other hand, sparse random matrices also attract much attention [26, 27, 28, 29, 30, 31, 32] since the sparsity can enable the computation of the matrix-product to be remarkably efficient and save the storage space in practice [33]. In particular, sparse binary-valued measurement matrices are commonly used in some applications, including group testing [34], DNA Microarrays [35], and single-pixel imaging [36]. Furthermore, many studies [29, 30, 37] showed that the sparse binary-valued measurement matrices are as “good” as the dense ones both in theory and in practice.
@@
In this paper, we focus on a type of sparse binary measurement matrix known as a sparse random combinatorial matrix. This matrix $\boldsymbol{A}\in\{0,1\}^{m\times n}$ is constructed such that each column is generated independently and uniformly at random from the set of all binary vectors of length $m$ that contain exactly $d$ ones, where $d\leq m/2$. In particular, sparse random combinatorial matrices have some favorable properties. It has been shown in [38, 39] that an $m\times m$ sparse random combinatorial matrix is nonsingular with probability 1o(1)1-o(1) for a sufficiently large $m$, if the degree $d$ of each column satisfies $(1+\gamma)\ln m\leq d\leq m/2$ for a constant $\gamma>0$. Now we answer the previous question. Yes, much redundancy of $\boldsymbol{A}$ can be removed if $\boldsymbol{x}$ is a signal defined in Definition 2. Actually, this observation can be traced back to the work of Khajehnejad et al. in [28]. They showed that the redundancy of a sparse matrix constructed by the expander theory can be eliminated if nonzero elements of $\boldsymbol{x}$ are non-negative. Our work is more general and takes the observation in [28] as a special case. Specifically, the sparse signal we consider is not limited to be non-negative, but a class of signals defined in Definition 2, including the Gaussian sparse signal. Furthermore, the considered measurement matrix $\boldsymbol{A}$ in this paper is more general than that constructed by the expander theory in [28]. The contributions of this paper are summarized as follows.
@@
  • We first define the confined set $\Gamma$ with $|\Gamma|\leq n$ and prove that the support of $\boldsymbol{x}$ defined in Definition 2 is a subset of $\Gamma$ with probability 1. To theoretically clarify the effectiveness of removing the redundancy of $\boldsymbol{A}$, we present the expectations of the sparsity of $\boldsymbol{y}$ and the size of $\Gamma$.
@@
  • We propose a variant of the OMP algorithm, referred to as the confined OMP algorithm, by introducing the confined set $\Gamma$ into the identification step. The possibly chosen column indices are strictly confined to the confined set $\Gamma$. We further analyze the complexities of the OMP and confined OMP algorithms. The analysis results show that the identification efficiency of confined OMP algorithm is at least $\frac{nKd-K}{|\Gamma|Kd-K+n}$ times that of OMP algorithm. Furthermore, the experimental results show that the confined OMP algorithm achieves a large reduction in complexity if $K\ll m$.
  • We develop a lower bound on the probability of exact recovery of $\boldsymbol{x}$ defined in Definition 2 using the confined OMP algorithm over a sparse random combinatorial matrix. Theoretical results show that, if the spark of $\boldsymbol{A}$ is greater than $K$, $m=2\mathrm{e}K\ln(n-K)$ measurements are sufficient to guarantee that the probability of recovering a $K$-sparse signal using the proposed algorithm is no lower than $1-\frac{1}{n-K}$. In the asymptotic regime where both $m$ and $n$ tend to infinity with $n=m^{\tau}~(\tau>1)$, the proposed algorithm can exactly recover signals with sparsity $K=o\left(\frac{m}{\ln m}\right)$ with probability 1.
@@
The paper is organized as follows. We define the confined set $\Gamma$ and present the proposed algorithm in Sec. II. In Sec. III, the expectations of the sparsity of $\boldsymbol{y}$ and the size of $\Gamma$ are investigated, and then the recovery performance analysis of the confined OMP algorithm is provided. Experimental results are presented in Sec. IV and Sec. V concludes the paper.
@@
Notation: we use boldface lowercase letters to denote column vectors and boldface uppercase letters to denote matrices. The support of $\boldsymbol{x}$ is denoted by $\Omega$ and the complement of $\Omega$ is $\Omega^{c}=[n]\backslash\Omega=\{i:i\in[n],i\notin\Omega\}$, where $[n]$ represents the set $\{1,2,\cdots,n\}$. Let $x_{i}$ and $\boldsymbol{A}_{j}$ be the $i$-th element of $\boldsymbol{x}$ and the $j$-th column of $\boldsymbol{A}$, respectively. We use $|\Gamma|$ to denote the cardinality of a set $\Gamma$ and $A_{i,j}$ to denote the element of $\boldsymbol{A}$ located at the $i$-th row and the $j$-th column. We denote by $\boldsymbol{x}_{\Lambda}$ the sub-vector of $\boldsymbol{x}$ that contains the entries of $\boldsymbol{x}$ indexed by the set $\Lambda$, and $\boldsymbol{A}_{\Lambda}$ the sub-matrix of $\boldsymbol{A}$ that contains the columns of $\boldsymbol{A}$ indexed by the set $\Lambda$. We use $\boldsymbol{A}^{T}$ to denote the transpose of $\boldsymbol{A}$ and $\mathbf{E}[X]$ to represent the expectation of $X$.
II Proposed Algorithm
@@
This section will first introduce the inherent properties of random combinatorial matrices and sparse signals, and then present the details of the proposed algorithm.
II-A The Confined Set
@@
For a real number $\epsilon>0$, let $\mathcal{E}=\{i:|y_{i}|\leq\epsilon\}$ where $y_{i}$ is the $i$-th element of $\boldsymbol{y}$. The definition of the confined set is as follows.
Definition 1 (The confined set)
@@
Theorem 1 (Lower bound on $\mathbb{P}\{\Omega\subseteq\Gamma\}$)
Suppose that in (1), $\boldsymbol{A}\in\{0,1\}^{m\times n}$ is a random combinatorial matrix with independent columns, where each column is chosen uniformly among the vectors with $d$ ones. Furthermore, the $K$ nonzero components of $\boldsymbol{x}$ are independent and identically distributed (i.i.d.), with the same cumulative distribution function (CDF) $F_{X}(x)=\mathbb{P}\{X\leq x\}$ and probability density function (PDF) $f_{X}(x)$. Then, the probability of $\{\Omega\subseteq\Gamma\}$ is lower bounded by
$\begin{split}&\mathbb{P}\{\Omega\subseteq\Gamma\}\geq\\ &1-m\sum_{\ell=1}^{K}\binom{K}{\ell}\left(\frac{d}{m}\right)^{\ell}\left(1-\frac{d}{m}\right)^{K-\ell}\Big(F_{X}^{*\ell}(\epsilon)-F_{X}^{*\ell}(-\epsilon)\Big),\end{split}$ (3)
@@
Proof:
See Appendix A. ∎
With Theorem 1, we have the following two Corollaries.
Corollary 1
If the PDF $f_{X}(x)$ of nonzero components of $\boldsymbol{x}$ is a continuous function, then the probability $\mathbb{P}\{\Omega\subseteq\Gamma\}\to 1$ as $\epsilon\to 0$.
Proof:
Obviously, $F_{X}^{*\ell}(x)$ is a continuous function provided that $f_{X}(x)$ is a continuous function. According to Theorem 1, we have $\underset{\epsilon\to 0}{\lim}\mathbb{P}\{\Omega\subseteq\Gamma\}=1$. ∎
Corollary 2
If the values of nonzero components of $\boldsymbol{x}$ share the same polarity, then the probability $\mathbb{P}\{\Omega\subseteq\Gamma\}\to 1$ as $\epsilon\to 0$.
The Corollaries indicate that by setting $\epsilon\to 0$, both Gaussian sparse signals and non-negative sparse signals can ensure $\mathbb{P}\{\Omega\subseteq\Gamma\}=1$. We call these signals confined signals in this paper. The definition of the confined signal is given as follows.
Definition 2 (The confined signal)
The confined signal is a $K$-sparse signal whose nonzero components are i.i.d. and, for $\ell=1,2,\cdots,K$, their CDFs satisfy $F_{X}^{*\ell}(\epsilon)-F_{X}^{*\ell}(-\epsilon)\to 0$ as $\epsilon\to 0$.
II-B Confined OMP Algorithm
The confined OMP is a modification of the OMP algorithm that introduces a confined set $\Gamma$. The key feature of the confined OMP algorithm is to introduce a confined set $\Gamma$ into the identification step such that the possibly chosen column indices are strictly confined to the confined set $\Gamma$. In the following, we assume that the signal $\boldsymbol{x}$ is a confined signal defined in Definition 2 and has exactly $K$ nonzero elements. The details of the confined OMP algorithm are summarized in Algorithm 2. The necessary explanations of the confined OMP algorithm are as follows.
[…]